# SKETCH: Semantic Key-Point Conditioning for Long-Horizon Vessel Trajectory Prediction

### v1

@@ 1 Introduction
More recently, Transformer-based models have shown promise due to their ability to model long-range dependencies via self-attention (Vaswani et al., 2017). Approaches such as TrAISformer enhance medium-term trajectory prediction and more effectively capture multimodal uncertainty (Nguyen and Fablet, 2024). Nevertheless, we observe that standard autoregressive Transformers still degrade severely in long-horizon settings. Notably, both TrAISformer and recent token-based GPT-style forecasters, such as TrackGPT (Stroh, 2024), convert continuous trajectory attributes into discrete tokens and learn embeddings over these discrete representations as the input to a Transformer backbone, which will prevent unseen embedded layer regions from being trained and will be fail when trajectories enter those regions. As uncertainty accumulates, predictions tend to collapse toward nearly horizontal trajectories. This observation suggests that increasing model capacity alone is insufficient for robust long-term trajectory forecasting.
@@ 3.2 Semantic Next-Key-Point Modeling
Definition 3.1 (Semantic Navigational Key Point (NKP)).
A Semantic Navigational Key Point (NKP) is defined as a latent variable that represents high-level navigational intent of a vessel, rather than a deterministic geometric waypoint. Specifically, an NKP corresponds to an equivalence class of future trajectories that share the same navigational semantics, such as passing through a specific channel, approaching a port region, or following a major shipping lane.
@@
To improve efficiency, we freeze Encoder 1 and MiniMind Model 1 after pretraining and only fine-tune a lightweight multi-layer perceptron (MLP) for NKP verification. In this way, the learned representation acts as a compact semantic embedding, capturing high-level navigational intent. In practice, we approximate marginalization over $Z$ using a MAP estimate, which performs robustly in our experiments.
During inference, an unknown trajectory is compared against a database of reference trajectories. For each reference trajectory whose similarity exceeds a predefined threshold, a vote is cast for its associated NKP. The NKP with the highest vote count is selected as the prediction, as summarized in Algorithm 1. This voting procedure can be interpreted as a non-parametric approximation of $p(Z\mid X)$ through nearest-neighbor evidence aggregation. The overall training and inference pipeline is shown in Fig. 2.
@@ 3.4 Model Architecture and Physical Representation
The derivation follows from integrating the velocity field under a spherical-Earth approximation and is presented in Appendix B. Fig. 3 clearly indicates that our assumption holds, with purple and green points representing prediction and ground truth, respectively.
Figure 3: from COG & SOG to latitude & longitude - dense. Purple and green points represent the next coordinates derived by the formula and ground truth, respectively. The fact that the purple and green dots are close enough indicates the practical validity of the formulation.
Table 1: Qualitative comparison of MP-LSTM, TrAISformer, and our model. They are trained on the same training split with the same procedure, and evaluated on the same test split.
Ours [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
MP-LSTM [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
trAISformer [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 2: Ablation Experiment for NKP Prediction. 6ch refers to the model with correct NKP, while 6ch w/ WD refers to the model with random NKP.
@@ 4 Evaluation Metrics
To comprehensively evaluate trajectory prediction performance at different granularities, we adopt three complementary metrics: Mean Squared Error of Position (MSEP), Mean Squared Error of Curvature (MSEC), Mean Fréchet Distance (MFD) and inference duration. Specifically, MSEP measures point-wise positional accuracy and reflects local coordinate-level errors. MSEC evaluates discrepancies in curvature profiles and characterizes the smoothness and local geometric consistency of predicted trajectories. In contrast, MFD captures curve-level similarity by assessing the global geometric alignment between predicted and ground-truth trajectories. The inference durations, in seconds, evaluate the model’s efficiency and include the time taken on the same test datasets at the same batch sizes. Together, these metrics provide a multi-scale evaluation framework that jointly accounts for local accuracy, trajectory smoothness, global shape consistency, and time efficiency. More datasets and implementation details were covered in Appendix Appendix C and Appendix D.
@@ 5.1 Metrics on Trajectory Prediction
Overall, our method achieves the best performance across all reported metrics. Specifically, our model attains the lowest MSEP, indicating more accurate point-wise predictions over the entire time horizon. At the curve level, our approach also achieves the smallest MFD, reflecting closer geometric and curvature-level alignment with the ground-truth trajectories. In addition, our model demonstrates a clear advantage in inference efficiency. As shown in Table 3, it requires significantly less inference time than both MP-LSTM and TrAISformer, making it more suitable for large-scale and real-time vessel trajectory prediction scenarios.
Table 3: Quantitative comparison of trajectory prediction performance between our method, MP-LSTM, and TrAISformer, evaluated using MSEP, MSEC, MFD, and inference time on the whole test dataset.
@@ 5.2 Generalization to Public AIS Dataset
Table 4: Metrics for trajectory prediction on the public AIS_Dataset (Xie et al., 2025).
@@ 5.3 Next Key Point Prediction Accuracy
As shown in Table 5, all our proposed models significantly outperform the Random Forest baseline (Zhang et al., 2020). Among them, pretrain-c achieves the highest closed-set accuracy (98.66%), while sft-c also performs strongly (95.44%). The open-set fine-tuned sft-o-s2 attains slightly lower accuracy (93.61%) but offers the advantage of seamless generalization to new NKPs through simple database extension, without retraining. This demonstrates both the effectiveness of the fine-tuning procedure and its potential for application to other evolving prediction tasks.
@@ 5.4 Effect of NKP Quality on Trajectory Prediction
We analyze the impact of Next Key Point (NKP) quality on downstream trajectory prediction in Table LABEL:tag:_stage_3_compare_with_approaches. The correct and wrong NKP settings define oracle upper and adversarial lower bounds, within which all models using predicted NKPs fall, indicating that NKP prediction errors induce only limited and acceptable degradation.
@@
Importantly, most NKP-conditioned models outperform the pure 4-channel baseline, confirming that incorporating NKPs as high-level intent provides critical global guidance beyond local motion cues, while incorrect NKPs cause severe trajectory distortion. Overall, these results demonstrate that NKP quality is crucial for achieving smooth and globally consistent long-horizon trajectory prediction.
@@ 5.5 Qualitative Results and Visualization
We qualitatively evaluate trajectory predictions against ground truth across different regions (Table 1). MP-LSTM (Gao et al., 2021), reimplemented using a training-set database, predicts destinations accurately but exhibits large deviations at support points. TrAISformer (Nguyen and Fablet, 2024), applied from the official repository, shows pauses and stagnant latitude. In contrast, our model accurately captures both straight and turning trajectories, maintaining correct timing and coordinates, which demonstrates robust long-horizon prediction in complex scenarios.
Table 2 highlights the effectiveness of the proposed NKP channel. Models with NKP consistently produce trajectories closer to the ground truth compared to those without, using the same random seed for fair comparison. The 6-channel row with an incorrect NKP input shows that the model effectively decouples NKP information: errors in NKP cause the predicted trajectory to twist, confirming the model’s sensitivity to NKP guidance while retaining robustness.
@@ A.5 Remarks on Stronger Monotonicity Notions
Appendix B Detailed Derivation of Next-Timestamp Coordinates
@@
Appendix C Dataset and Data Preprocessing
C.1 Data Description and Coverage for Private Dataset
@@
C.2 Data Preprocessing Pipeline
@@
C.3 Stage 2 Database Sampling Strategy
@@
C.4 Data Description and Coverage for Public Dataset
@@
Appendix D Implementation Details
D.1 Model Implementation
@@ D.2 Training Details
Appendix E Broader Impact

### v2

@@ 1 Introduction
More recently, Transformer-based models have shown promise due to their ability to model long-range dependencies via self-attention (Vaswani et al., 2017). Approaches such as TrAISformer enhance medium-term trajectory prediction and more effectively capture multimodal uncertainty (Nguyen and Fablet, 2024). Nevertheless, we observe that standard autoregressive Transformers still degrade severely in long-horizon settings. Notably, both TrAISformer (Nguyen and Fablet, 2024) and recent token-based GPT-style forecasters, such as TrackGPT (Stroh, 2024), convert continuous trajectory attributes into discrete tokens and learn embeddings over these discrete representations as the input to a Transformer backbone. Due to the discretization of continuous geographic space, token-based trajectory models are trained only on a sparse subset of spatial tokens. When deployed in previously unseen regions, the model must extrapolate beyond the empirical support of the training distribution, often leading to unstable or incoherent predictions. As uncertainty accumulates, predictions tend to collapse toward nearly horizontal trajectories. This observation suggests that increasing model capacity alone is insufficient for robust long-term trajectory forecasting.
@@ 3.2 Semantic Next-Key-Point Modeling
Definition 3.1 (Semantic Next Key Point (NKP)).
A Semantic Next Key Point (NKP) is a latent semantic variable that represents the anticipated next navigational anchor of a vessel, inferred from its observed trajectory history. Unlike a deterministic geometric waypoint, an NKP encodes high-level future intent by characterizing an equivalence class of plausible future trajectories that share consistent navigational semantics, such as heading toward a port region, passing through a specific strait, or following a major shipping lane. NKPs are not the terminal points, and by introducing NKPs, an extremely long trajectory can be decomposed into several shorter segments, each governed by a coherent semantic intent, enabling hierarchical long-horizon prediction.
Remark. We use the term semantic to emphasize that an NKP represents an intent-level equivalence class of feasible futures rather than a precise geometric waypoint, although in practice we obtain NKP supervision from spatial annotations (e.g., port/strait intersections) for training and evaluation.
@@
To improve efficiency, we freeze Encoder 1 and MiniMind Model 1 after pretraining and only fine-tune a lightweight multi-layer perceptron (MLP) for NKP verification. In this way, the learned representation acts as a compact semantic embedding, capturing high-level navigational intent. In practice, we approximate marginalization over $Z$ using a MAP estimate, which performs robustly in our experiments. When $P(Z\mid X)$ is sharply peaked, marginalization can be well approximated by conditioning on the MAP NKP.
During inference, an unknown trajectory is compared against a database of reference trajectories. The reference database contains only training trajectories for both the public and private test datasets. For each reference trajectory whose similarity exceeds a predefined threshold, a vote is cast for its associated NKP. The NKP with the highest vote count is selected as the prediction, as summarized in Algorithm 1. This voting procedure can be interpreted as a nonparametric approximation to $p(Z\mid X)$ via nearest-neighbor evidence aggregation. The overall training and inference pipeline is shown in Fig. 2.
@@ 3.4 Model Architecture and Physical Representation
This formulation operates in a locally Euclidean space and avoids repeated spherical projections, thereby improving numerical stability and preventing systematic error amplification during long-horizon autoregressive rollout. The derivation follows from integrating the velocity field under a spherical-Earth approximation and is presented in Appendix C, along with a numerical stability analysis in Appendix D.
To empirically assess the numerical consistency of the proposed coordinate update, we perform a one-step displacement check on real AIS trajectories. For each trajectory segment of length $288$, the next-step position is computed from the current SOG/COG and compared against the ground-truth next coordinate. Across the evaluated trajectories, the resulting mean squared error is on the order of $10^{-9}$, indicating numerical consistency up to machine precision.
@@ 4 Evaluation Metrics
To comprehensively evaluate trajectory prediction performance at different granularities, we adopt three complementary metrics: Mean Squared Error of Position (MSEP), Mean Squared Error of Curvature (MSEC), Mean Fréchet Distance (MFD) and inference duration. Specifically, MSEP measures point-wise positional accuracy and reflects local coordinate-level errors. MSEC evaluates discrepancies in curvature profiles and characterizes the smoothness and local geometric consistency of predicted trajectories. In contrast, MFD captures curve-level similarity by assessing the global geometric alignment between predicted and ground-truth trajectories. The inference durations, in seconds, evaluate the model’s efficiency and include the time taken on the same test datasets at the same batch sizes. Together, these metrics provide a multi-scale evaluation framework that jointly accounts for local accuracy, trajectory smoothness, global shape consistency, and time efficiency. More datasets and implementation details were covered in Appendix Appendix E and Appendix F.
Table 1: Qualitative comparison of MP-LSTM, TrAISformer, and our model. They are trained on the same training split with the same procedure, and evaluated on the same test split. Blue, green, red, orange, and purple trajectories refer to input, ground truth, our prediction, prediction from MP-LSTM (Gao et al., 2021), and prediction from TrAISformer (Nguyen and Fablet, 2024), respectively. Additional qualitative examples are provided in Appendix B.
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@ 5.1 Metrics on Trajectory Prediction
Overall, our method achieves the best performance across all reported metrics. Specifically, our model attains the lowest MSEP, indicating more accurate point-wise predictions over the entire time horizon. At the curve level, our approach also achieves the smallest MFD, reflecting closer geometric and curvature-level alignment with the ground-truth trajectories.
In addition, our model demonstrates a clear advantage in inference efficiency. As shown in Table 2, it requires significantly less inference time than both MP-LSTM and TrAISformer, making it more suitable for large-scale and real-time vessel trajectory prediction scenarios.
Table 2: Quantitative comparison of trajectory prediction performance between our method, MP-LSTM, and TrAISformer, evaluated using MSEP, MSEC, MFD, and inference time on the whole test dataset. The unit of time is second, and was tested on an NVIDIA RTX 4090 for the whole test dataset with 2513 samples, and a batch size of 128.
@@ 5.2 Generalization to Public AIS Dataset
Table 3: Metrics for trajectory prediction on the public AIS_Dataset (Xie et al., 2025). All the metrics are the same as Table 2 with a total of 560 samples.
@@ 5.3 Next Key Point Prediction Accuracy
As shown in Table 4, all our proposed models significantly outperform the Random Forest baseline (Zhang et al., 2020). Among them, pretrain-c achieves the highest closed-set accuracy (98.98%), while sft-c also performs strongly (95.44%). The open-set fine-tuned sft-o-s2 attains slightly lower accuracy (93.61%) but offers the advantage of seamless generalization to new NKPs through simple database extension, without retraining. This demonstrates both the effectiveness of the fine-tuning procedure and its potential for application to other evolving prediction tasks.
@@ 5.4 Effect of NKP Quality on Trajectory Prediction
We analyze the impact of Next Key Point (NKP) quality on the prediction of downstream trajectory in Table 5. The correct and wrong NKP settings define oracle upper and adversarial lower bounds, within which all models using predicted NKPs fall, indicating that NKP prediction errors induce only limited and acceptable degradation.
@@
Table 6: Ablation Experiment for NKP Prediction. 4ch refers to the model without NKP information, 6ch refers to the model with correct NKP, while 6ch w/ WD refers to the model with random but wrong NKP.
@@
Importantly, most NKP-conditioned models outperform the pure 4-channel baseline, confirming that incorporating NKPs as high-level intent provides critical global guidance beyond local motion cues, while incorrect NKPs cause severe trajectory distortion. Overall, these results demonstrate that the quality of NKP is crucial to achieving smooth and globally consistent long-horizon trajectory prediction.
@@ 5.5 Qualitative Results and Visualization
We qualitatively evaluate trajectory predictions against the ground truth across different regions (Table 1). MP-LSTM (Gao et al., 2021) predicts trajectories using only two anchor points (support and destination). While the support point is generally more accurate, the destination is often less reliable, and an error in either anchor can induce a significant global deviation in the reconstructed trajectory. TrAISformer (Nguyen and Fablet, 2024), applied from the official repository, shows pauses and stagnant latitude. In contrast, our model accurately captures both straight and turning trajectories, maintaining correct timing and coordinates, which demonstrates robust long-horizon prediction in complex scenarios.
Table 6 highlights the effectiveness of the proposed NKP channel. Models with NKP consistently produce trajectories closer to the ground truth compared to those without it, using the same random seed for fair comparison. The 6-channel row with an incorrect NKP input shows that the model effectively decouples NKP information: errors in NKP cause the predicted trajectory to twist, confirming the model’s sensitivity to NKP guidance while retaining robustness.
@@ A.5 Remarks on Stronger Monotonicity Notions
Appendix B Additional Qualitative Results
In this appendix, we provide additional qualitative comparisons between MP-LSTM, TrAISformer, and our method on diverse navigation scenarios.
Table 7: Additional qualitative trajectory prediction examples.
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Appendix C Detailed Derivation of Next-Timestamp Coordinates
@@
Appendix D Numerical Stability of Autoregressive Rollout
We analyze the numerical behavior of autoregressive trajectory rollout under different coordinate update formulations.
Let $s_{t}\in\mathbb{R}^{d}$ denote the true vessel state at time $t$ and $\hat{s}_{t}$ the predicted state, with prediction error $e_{t}=\hat{s}_{t}-s_{t}$. In autoregressive forecasting, predictions are propagated according to
$\hat{s}_{t+1}=f(\hat{s}_{t}),$ (36)
which induces the first-order error dynamics
$e_{t+1}\approx J_{f}(s_{t})\,e_{t},$ (37)
where $J_{f}$ denotes the Jacobian of the state transition function evaluated at $s_{t}$. Long-horizon numerical stability therefore depends critically on the spectral properties of $J_{f}$.
D.1 Spherical Geometry–Based Updates
In spherical or geodesic formulations, state updates are typically implemented as a composition of coordinate projection and motion integration:
$f_{\mathrm{sph}}=\Pi^{-1}\circ G\circ\Pi,$ (38)
where $\Pi$ maps geographic coordinates to a local tangent plane, and $G$ performs motion updates in that plane.
The Jacobians of $\Pi$ and $\Pi^{-1}$ depend nonlinearly on latitude and heading, introducing location-dependent scaling factors. The resulting error propagation becomes
$e_{t+1}\approx J_{\Pi^{-1}}\,J_{G}\,J_{\Pi}\,e_{t}.$ (39)
Even if $J_{G}$ is well-conditioned, repeated application of $J_{\Pi}$ and $J_{\Pi^{-1}}$ across time steps leads to multiplicative error accumulation. As a result, small numerical inaccuracies can be systematically amplified during long-horizon rollout, especially at high latitudes or under frequent heading changes.
D.2 Locally Euclidean Update (Ours)
In contrast, our formulation performs state updates directly in a locally Euclidean coordinate system consistent with AIS-reported SOG and COG. Under the local linear motion assumption, the update function is smooth and approximately affine over short time intervals:
$f_{\mathrm{ours}}(s_{t})=s_{t}+\Delta t\cdot v(s_{t}),$ (40)
where $v(\cdot)$ denotes the velocity field induced by SOG and COG.
The corresponding Jacobian satisfies
$J_{f}\approx I+\mathcal{O}(\Delta t),$ (41)
where $I$ is the identity matrix. Consequently, prediction errors accumulate approximately additively rather than multiplicatively:
$\|e_{T}\|\leq\|e_{0}\|+\mathcal{O}(T\Delta t).$ (42)
This behavior contrasts sharply with spherical formulations, where the effective Jacobian deviates from identity due to repeated nonlinear projections. Therefore, our locally Euclidean update is numerically more stable under long-horizon autoregressive rollout and prevents systematic drift caused by geometric distortion.
Appendix E Dataset and Data Preprocessing
E.1 Data Description and Coverage for Private Dataset
@@
E.2 Data Preprocessing Pipeline
@@
E.3 Stage 2 Database Sampling Strategy
@@
E.4 Data Description and Coverage for Public Dataset
@@
Appendix F Implementation Details
F.1 Model Implementation
@@ F.2 Training Details
Appendix G Broader Impact